Find the equation of the line parallel passing through 3x+y=7. x=2;y=1 3/4
Answer by nerdybill(5404)

Find the equation of the line parallel passing through 3x+y=7. x=2;y=1 3/4
.
Begin by finding the slope of:
3x+y=7
Do this by manipulating the above until it is in the form of:
y = mx + b
.
3x+y=7
y = -3x + 7
Here, m=-3 (slope)
.
Now, we KNOW that parallel lines have to have identical slopes. Therefore, our new equation will have a slope of -3.
.
The problem also wanted the equation to pass through:
x=2;y=1 3/4
or
(x,y) = (2,1.75)
.
Plug the above and the slope into the "slope-intercept" form of the line and solve for b:
y = mx + b
1.75 = -3(2) + b
1.75 = -6 + b
7.75 = b
.
Therefore, the equation of the line is:
y = -3x + 7.75